Drops are falling regularly from a water tap at a height of 9 metre from the ground. The 4th drop is about to fall from the tap when the 1st hits the ground. Find the distance between 2nd and 3rd drop?

Grade:11

1 Answers

Sandeep Bhamoo
5 years ago
Let the drops be falling at an intetval of "t" then displacement (s=at²) of fitst drop is =a(3t)²=9k =9 ,hence k=1, now for second drop s=4k=4 and for third s=k=1m distance b/w second and third dtop is = 4-1=3. Please do approve my answer if correct